What is y = (x + 3)2 + (x + 4)2 rewritten in vertex form?

Answer by fractalier(6550)

y = (x+3)^2 + (x+4)^2
y = x^2 + 6x + 9 + x^2 + 8x + 16
y = 2x^2 + 14x + 25
y = 2(x^2 + 7x) + 25
Now complete the square...
Half of 7, squared, is 49/4...so we add that and subtract it giving us
y = 2(x^2 + 7x + 49/4) + 25 - 49/2
y = 2(x + 7/2)^2 + 1/2